Fermi National Laboratory


Accelerator Update

Wednesday March 24
The day shift began with the Antiproton Source (Pbar) stacking, with Recycler experts conducting studies, with safety system techs testing the Meson safety system, and with feeder 23 repairs at the Master Substation.

At 9:37 AM, Operations reported seeing a power glitch that caused problems in Linac, Booster, and Pbar. CHL also saw the glitch, but suffered no harm. The Pbar stack remained intact. The High Voltage expert reported that ComEd had a transformer fail in Plano, Illinois.

Tevatron (TeV) experts reported at 9:51 AM that the cryo system at TeV sector B1 would have to warm up in order to do some alignments.

Pbar and the Main Injector (MI) experts prepared their machines at 10:27 AM for a shot setup to verify systems.

Pbar resumed stacking at 1:40 PM.

At 3:57 PM, Operations began sending beam to Switchyard (SY) in an attempt to get the beam to the SY dump.

Both the Neutron Therapy Facility (NTF) and Pbar began study periods at 4 PM.

At 7:23 PM, a SY expert got beam to the SY dump.

Alignment techs reported at 9:09 PM that they had completed aligning the TeV components at the Collider Detector at Fermilab (CDF). Riggers began replacing the steel blocks. Cryo system cooldown will began when the riggers finish.

NTF completed its studies at 9:30 PM.

At 10:23 PM, operators finished interlocking the Meson enclosures.

Thursday March 25
The midnight shift began with operators interlocking CDF, with Pbar and Recycler conducting studies, and with Operations preparing for the TeV turn on.

Pbar studies ended at 12:15 AM.

After a lot of work by TeV experts, techs, and operators Operations established a 980 GeV TeV ramp at 4:17 AM.

At 6:16 AM, Operations encountered some TeV problems while conducting a dry squeeze. Operators changed out a control card (several times) to one that had new firmware. At 7:07 AM, Operators put the TeV into another dry squeeze.

Extractor voltage for the H- Source began to wander at 7:20 AM. An expert was here and investigated. (Later in the day the I- Source was removed for repair; it will be reinstalled in about two weeks.)

By 8 AM, Pbar, TeV, and Recycler were all conducting studies.

Due to a Pbar problem that required a tunnel access, Pbar and Recycler went into shot setup at 10:14 AM. After a slow tune-up, Recycler began stashing antiprotons at 12:46 PM. MI also used some of the antiprotons for study purposes. Recycler finished stashing at 2:20 PM.

Operators had the Pbar tunnels ready to access by 2:35 PM. The techs will investigate Debuncher cooling amplifier problems.

Booster collimator studies ended around 4:39 PM.

Operation began preparing MiniBooNE and SY for beam at 6:07 PM.

MI suffered a vacuum burst in its RF sector at 10:07 PM. Everything reset okay.

Experts finished their TeV ramp studies at 11:37 PM.

Friday March 26
The midnight shift began with Pbar stacking, with Operations monitoring the Recycler stash, with TeV studies, and with MiniBooNE taking beam.

Operations reported at 3:44 AM that a Booster RF station (BRF16) had tripped off three times in a short period of time. Operators bypassed the station.

Vacuum techs accessed the SY F1 enclosure at 5:30 AM to work on the Septa vacuum. They completed their work an hour and a half later.

TeV studies ended at 6:48 AM.

At 7:02 AM, operators had the D-zero (D0) collision hall ready for access and CDF ready by 7:33 AM.

Today and the Weekend Plans
The TeV experts will continue with studies and resolving problems. Their hope is to establish a store sometime during the Saturday day shift. All other machines will continue with necessary studies.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 3/26/2004   email Fermilab